OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was 3/8 the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television? 7.5 inches by 15 inches 28 inches by 56 inches 32 inches by 64 inches 4.5 inches by 9 inches

OpenStudy (anonymous):

is the answer B?

OpenStudy (anonymous):

Nope. The generic version was based off the brand name and was 3/8 the size of the brand name. Generic = (3/8)Brand Multiply both sides by (8/3) to eliminate the fraction on the right and you get: (8/3)Generic = Brand (8/3)(12) = ? (8/3)(24) = ?

OpenStudy (anonymous):

The answer is C?

OpenStudy (anonymous):

Sidenote: I think you were on the right track, but may have typed in (7/3) instead of (8/3), because: 12(7/3) = 28 (24(7/3) = 56

OpenStudy (anonymous):

Yep, C.